OpenStudy (dugalde):

use the half-angle formula to evaluate tan(17pi/12

OpenStudy (anonymous):

Do you know the half angle formula for tan?

OpenStudy (dugalde):

no

OpenStudy (anonymous):

\[\tan \frac{ \theta }{ 2 }=\frac{ \tan \theta }{ 1+\sqrt{1+ \tan^2 \theta} }\]

OpenStudy (anonymous):

In your equation,\[\frac{ \theta }{ 2 }=\frac{ 17\pi }{ 12 }\] so solve for Θ